# Fractality of open clusters in singles, pairs, and groups

### v1

@@ 5 Conclusions
These results demonstrate that both the intrinsic properties of clusters and their environmental context strongly influence their evolution. The higher fractality observed in clusters within pairs and groups indicates that their structural development remains closely tied to the initial substructure and dynamics of their parental molecular clouds. While more massive clusters tend to evolve toward centrally concentrated, radially symmetric configurations, less massive clusters are able to retain their fractal substructure for longer periods. Overall, our findings emphasize that open clusters do not evolve in isolation: single clusters exhibit clear signs of advanced dynamical relaxation, whereas clusters in pairs and groups preserve structural imprints of their formation environment, highlighting the crucial role of environmental interactions in shaping their long-term evolution.

### v3

@@ 6 Conclusion
These results demonstrate that both the intrinsic properties of clusters and their environmental context strongly influence their evolution. The higher fractality observed in clusters within pairs and groups indicates that their structural development remains closely tied to the initial substructure and dynamics of their parental molecular clouds. While more massive clusters tend to evolve toward centrally concentrated, radially symmetric configurations, less massive clusters are able to retain their fractal substructure for longer periods. Overall, our findings emphasize that OCs do not evolve in isolation: single clusters exhibit clear signs of advanced dynamical relaxation, whereas clusters in pairs and groups preserve structural imprints of their formation environment, highlighting the crucial role of environmental interactions in shaping their long-term evolution.